OpenStudy (anonymous):

(ANSWER PLEASE) Let (sin x)(cos x) = .35 and sin x + cos x = 1.30. Which is (csc x) + (sec x)? A. 3.71 B. 2.17 C. 3.77 D. 2.71

OpenStudy (loser66):

3.71

OpenStudy (loser66):

\[\frac{sinx}{0.35}=\frac{1}{cosx}=sec x\] \[\frac{cos x}{0.35}=\frac{1}{sinx}=csx x\] plus them together to get the answer, that's it

OpenStudy (campbell_st):

think of it this way \[\csc(x) + \sec(x) = \frac{1}{\sin(x)} + \frac{1}{\cos(x)} = \frac{\sin(x) + \cos(x)}{\sin(x)\cos(x)}\] then its numeric substitution